# Diederich–Fornæss index and global regularity in the $\overline{\partial}$–Neumann problem: domains with comparable Levi eigenvalues

### v4

@@ 4. Proof of Theorem 1
The first three error terms in the last line are the same as in (40) and so are acceptable, as is the last. For fourth one, we use the subelliptic estimate $\|(N_{\delta,q})_{N}\|_{k}^{2}\lesssim\|\overline{\partial}N_{\delta,q}u\|_{k-1}^{2}+\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}+\|N_{\delta,q}u\|_{k-1}^{2}$ ([36], estimate (2.96)). Using (28) again and the induction hypothesis shows that this term is also acceptable.
@@
For $j<n$, the contributions coming from $\overline{L_{j}}(\rho g_{\eta})$ and $L_{j}(\rho g_{\eta})$, respectively, are $\mathcal{O}_{\eta}(\|\rho(X_{\eta}-\overline{X_{\eta}})^{k}\chi N_{\delta,q}u\|_{k}$ and are again acceptable ($\rho$ is a subelliptic multiplier and (28)). Because $\alpha_{\eta}$ is real, $|\alpha_{\eta}(\overline{L_{j}})|^{2}=|\alpha_{\eta}(L_{j})|^{2}$, and we only have to estimate $\int_{\Omega}|\alpha_{\eta}(L_{j})|^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}=\int_{\Omega}|\alpha_{\eta}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}L_{j})|^{2}$. We do this via (14) in Proposition 1, applied to the $(0,1)$–form $(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}\overline{\omega_{j}}$. Then $L_{u}=((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}L_{j}$, and the resulting estimate is
(42) $\int_{\Omega}|\alpha_{\eta}(L_{j})|^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}\overline{\omega_{j}})\|^{2}+\|\overline{\partial}^{*}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}\overline{\omega_{j}})\|^{2}\right)\\ +C_{\eta}\|(X_{\eta}-\overline{X_{\eta}})^{k}N_{\delta,q}u\|_{-1}^{2}\;.$
We have again omitted the factor $\chi$ from the error term; commuting $\chi$ to the front gives a term that is of the order of $C_{\eta}\|N_{\delta,q}u\|_{k-1}$. These terms are acceptable; this is clear for the $\|\cdot\|_{k-1}$ term (also using (28)). For the others, it follows from interior elliptic regularity of $\Box_{\delta,q}$, with constants that are uniform in $\delta$. Once $\chi$ is first, the terms are bounded by the norms over all of $\Omega$.
The $\overline{\partial}$ and $\overline{\partial}^{*}$ terms in the second line of (42) are dominated (uniformly in $\eta$ and $\delta$) by
(43) $\sum_{s=1}^{n}\|\overline{L_{s}}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J})\|^{2}+\|L_{j}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J})\|^{2}+\|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)\|^{2}\\ \lesssim\|\overline{\partial}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u))\|^{2}+\|\overline{\partial}^{*}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u))\|^{2}\;;$
we have used the basic $L^{2}$ estimate for the first term, maximal estimates for the second (see (3); note again that $q_{0}\leq q$, and Lemma 2), and the basic $L^{2}$ estimate for the third.
@@
There remain the terms in (40) and (41) where $j=n$. In (40), we have that $|\overline{L_{n}}(\rho g_{\eta})-\alpha_{\eta}(\overline{L_{n}})|^{2}=0$ on the boundary, due to the choice of $g_{\eta}$. The contribution from this term is thus $\mathcal{O}_{\eta}(\|\rho(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)\|^{2})$. Again because $\rho$ is a subelliptic multiplier and (28), this term is acceptable. In (41), the term with $j=n$ occurs only in a sum over $J$ with $n\in J$. That is, $(\chi N_{\delta,q}u)_{J}$ is a component of the normal part $(\chi N_{\delta,q}u)_{N}$ of $\chi N_{\delta,q}u$ and so is subelliptic; we conclude as above that it is acceptable.
@@
The left hand side of (45) is what is needed to estimate the sum of the left hand sides of (31), (32), and (33). Thus

### v5

@@ 4. Proof of Theorem 1
The first three error terms in the last line are the same as in (39) and so are acceptable, as is the last. For the fourth one, we use the subelliptic estimate $\|(N_{\delta,q})_{N}\|_{k}^{2}\lesssim\|\overline{\partial}N_{\delta,q}u\|_{k-1}^{2}+\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}+\|N_{\delta,q}u\|_{k-1}^{2}$ ([36], estimate (2.96)). Using (28) again and the induction hypothesis shows that this term is also acceptable.
@@
For $j<n$, the contributions in (39) and (40) coming from $\overline{L_{j}}(\rho_{\eta}g_{\eta})$ and $L_{j}(\rho_{\eta}g_{\eta})$, respectively, are $\mathcal{O}_{\eta}(\|\rho(X_{\eta}-\overline{X_{\eta}})^{k}\chi N_{\delta,q}u\|_{k}$ and are again acceptable ($\rho_{\eta}$ is a subelliptic multiplier and (28)). For the term in (39) where $j=n$, we have that $\big{|}\overline{L_{n}^{\rho_{\eta}}}(\rho_{\eta}g_{\eta})-\alpha_{\eta}(\overline{L_{n}^{\rho_{\eta}}})|^{2}=0$ on the boundary, due to the choice of $g_{\eta}$ (note that $\overline{L_{n}^{\rho_{\eta}}}(\rho_{\eta}g_{\eta})=g_{\eta}$ on the boundary). The contribution from this term is thus $\mathcal{O}_{\eta}(\|\rho(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)\|^{2})$. Again because $\rho_{\eta}$ is a subelliptic multiplier and (28), this term is acceptable. In (40), $L_{n}(\rho_{\eta}g_{\eta})$ multiplies a coefficient of the normal component of $(X_{\eta}-\overline{X_{\eta}})\chi N_{\delta,q}u$. Taking the normal component acts like a subelliptic multiplier ([36], estimate (2.96)), so this term is also benign.
We now discuss the remaining terms in the second lines of (39) and (40), respectively. We want to apply Proposition 1. In order to do so, observe that it does not matter whether we form the vector fields $L^{J}_{u}$ in Euclidean coordinates or with coordinates in a special boundary frame: one set is obtained from the other via linear combinations (with smooth coefficients) when the from $u$ is fixed. Therefore, the validity of (14) is not affected. We start with (39); recall that for $j<n$, the terms containing $g_{\eta}$ have been taken care of, as has the term where $j=n$. Also, if $n\in J$, we can use the subelliptic estimate for the normal component of a form once more; these terms are thus acceptable. Because $\alpha_{\eta}$ is real, $|\alpha_{\eta}(\overline{L_{j}})|^{2}=|\alpha_{\eta}(L_{j})|^{2}$, and we only have to estimate $\int_{\Omega}|\alpha_{\eta}(L_{j})|^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}=\int_{\Omega}|\alpha_{\eta}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}L_{j})|^{2}$. Because $j\notin J$, $L_{\overline{\omega_{j}}\wedge\overline{\omega_{J}}}^{J}=\sum_{s=1}^{n}(\overline{\omega_{j}}\wedge\overline{\omega_{J}})_{sJ}L_{s}=L_{j}$ (the sum is only over $s\notin J$), and so $L^{J}_{\big{(}(X_{\eta}-\overline{X_{\eta}})^{k}\chi N_{\delta,q}u\big{)}\overline{\omega_{j}}\wedge\overline{\omega_{J}}}=\big{(}(X_{\eta}-\overline{X_{\eta}})^{k}\chi N_{\delta,q}u\big{)}_{J}L_{j}$. Proposition 1 applies to the $(0,q+1)$–form $\big{(}(X_{\eta}-\overline{X_{\eta}})^{k}\chi N_{\delta,q}u\big{)}_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J}})$; note that because $j<n$ and $n\notin J$, this form is in the domain of $\overline{\partial}^{*}$. The resulting estimate is
(41) $\sideset{}{{}^{\prime}}{\sum}_{J}\sum_{j\notin J,j<n}\int_{\Omega}|\alpha_{\eta}(L_{j})|^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}\\ \lesssim(1-\eta)\sideset{}{{}^{\prime}}{\sum}_{J}\sum_{j\notin J,j<n}\left(\|\overline{\partial}\big{(}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J})(\overline{\omega_{j}}\wedge\overline{\omega_{J}})\big{)}\|^{2}\right.\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left.\|\overline{\partial}^{*}\big{(}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J})(\overline{\omega_{j}}\wedge\overline{\omega_{J}})\big{)}\|^{2}\right)\\ +C_{\eta}\|(X_{\eta}-\overline{X_{\eta}})^{k}N_{\delta,q}u\|_{-1}^{2}\;.\;\;\;$
We have again omitted the factor $\chi$ from the error term; commuting $\chi$ to the front gives an error term that is of the order of $C_{\eta}\|N_{\delta,q}u\|_{k-1}$. These terms are acceptable. Once $\chi$ is first, the terms are bounded by the norms over all of $\Omega$. The $\overline{\partial}$ and $\overline{\partial}^{*}$ terms in the second and third lines of (41) are dominated (uniformly in $\eta$ and $\delta$) by
(42) $\sum_{s=1}^{n}\|\overline{L_{s}}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J})\|^{2}+\sum_{j<n}\|L_{j}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J})\|^{2}+\|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)\|^{2}\\ \lesssim\|\overline{\partial}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u))\|^{2}+\|\overline{\partial}^{*}((X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u))\|^{2}\;;$
we have used Lemma 1 for the first term, maximal estimates for the second (see (3); note again that $q_{0}\leq q$, and Lemma 2), and the basic $L^{2}$ estimate for the third. There is no $L_{n}$–term because the normal component of $\overline{\omega_{j}}\wedge\overline{\omega_{J}}$ is zero (as $j<n$, $n\notin J$).
In (40), the integrand equals $\big{|}\alpha_{\eta}\big{(}\sum_{j}(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{jS}L_{j}\big{)}\big{|}^{2}$ (again omitting the contribution from the $g_{\eta}$–term). The vector field inside $\alpha_{\eta}(\cdot)$ equals $L^{S}_{(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)}$. The form $(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)$ is in the domain of $\overline{\partial}^{*}$, and Proposition 1 applies and gives
(43) $\sideset{}{{}^{\prime}}{\sum}_{S}\int_{\Omega}\big{|}\sum_{j<n}\alpha_{\eta}(L_{j})(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{jS}\big{|}^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}\big{(}(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)\big{)}\|^{2}+\|\overline{\partial}^{*}\big{(}(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)\big{)}\|^{2}\right)\\ +C_{\eta}\|N_{\delta,q}u\|_{k-1}^{2}\;.$
@@
The left hand side of (45) is what is needed to estimate the sum of the left hand sides of (31), (32), and (33) (see the discussion form (34) to (38)). Thus